# Sampling as Bandits: Evaluation-Efficient Design for Black-Box Densities

### v1

@@ Abstract
We introduce bandit importance sampling (BIS), a new class of importance sampling methods designed for settings where the target density is expensive to evaluate. In contrast to adaptive importance sampling, which optimises a proposal distribution, BIS directly designs the samples through a sequential strategy that combines space-filling designs with multi-armed bandits. Our method leverages Gaussian process surrogates to guide sample selection, enabling efficient exploration of the parameter space with minimal target evaluations. We establish theoretical guarantees on convergence and demonstrate the effectiveness of the method across a broad range of sampling tasks. BIS delivers accurate approximations with fewer target evaluations, outperforming competing approaches across multimodal, heavy-tailed distributions, and real-world applications to Bayesian inference of computationally expensive models.
@@ 1 Introduction
Recent advances in machine learning and data science have broadened the scope of statistical analysis to encompass highly complex phenomena (Chen et al., 2010). The precise statistical description of complex phenomena often relies on highly-structured, black-box models that are computationally expensive to evaluate. Bayesian inference offers coherent probabilistic reasoning for deducing appropriate models from observations and prior knowledge. Yet, Bayesian inference can be computationally impractical for black-box models. Many cases emerge in, e.g., computational physics, biology, and geostatistics, where the model likelihoods depend on dynamical equations that need to be numerically solved (Beaumont et al., 2002; Girolami, 2008; Marin et al., 2012; Warne et al., 2020) or components involving exceedingly-large matrices and summations (Rue and Held, 2005; Goodreau et al., 2009; Besag, 1986).
@@
The chief concern in Bayesian inference is to sample from a posterior density of the model parameter, proportional to the model likelihood and a prior density. By design, standard sampling methods, such as Markov-Chain Monte Carlo (MCMC; Brooks et al., 2011), require a large number of evaluations of the target density for different parameter values. The computational cost quickly exceeds a desirable budget if the target density is expensive to evaluate. Thus, from a practical viewpoint, there is a need for novel sampling methods designed for black-box densities, which require only a small number of target evaluations.
@@
Importance sampling—originally introduced as a variance reduction technique in Monte Carlo estimation (Hammersley and Morton, 1954)—is an appealing approach in this context. The target density is approximated by a weighted average of samples drawn from an arbitrary proposal density (Tokdar and Kass, 2010). Given samples from the proposal density, evaluations of the target density occur only up to the number of the samples. However, the design of the samples is a determinant factor in the approximation accuracy of importance sampling. Adaptive importance sampling (AIS) is arguably the most common approach to the effective design of the samples, which optimises the proposal density so that it approximates the target density (e.g. Bugallo et al., 2017; Cotter et al., 2019, 2020). Yet, AIS can be easily impractical for black-box densities, as it reintroduces the need for intensive evaluation of the target density to optimise the proposal density.
In machine learning, Bayesian optimisation (BO; Mockus, 1989) has gained immense attention for problems involving black-box objective functions that are expensive to evaluate (Shahriari et al., 2016). The key idea is to replace the black-box objective with a computationally-cheap surrogate model constructed from the record of objective values evaluated in the past. Gaussian process (GP; Rasmussen and Williams, 2005) is a particularly successful choice in surrogate modelling (Gramacy, 2020). It not only provides predictions based on interpolation of the observed values, but also estimates the uncertainty in them arising from the non-identifiability of the true function given a finite amount of data. BO iteratively selects points that maximise a specified acquisition criterion based on the GP surrogate model. BO—sometimes referred to as GP bandits—can be viewed as a bandit problem with infinitely many arms (Bull, 2011). Hence, the framework of BO has been leveraged for experimental designs and active learning problems (Srinivas et al., 2010).
@@
This work establishes bandit importance sampling (BIS), a novel class of importance sampling that provides adaptively-designed samples. BIS retains the original advantage of importance sampling in that the target density is evaluated only up to the number of samples. It drastically reduces the number of samples compared to the level that standard MCMC or importance sampling typically requires for accurate approximation of the target density. The approach of BIS is to combine importance sampling with multi-armed bandits. AIS optimises the proposal density and draws samples randomly. In contrast, BIS optimises over the samples directly, sequentially selecting them from a pool of candidates based on a bandit strategy. The bandit strategy of BIS capitalises on GP surrogate modelling of the target density. While experimental designs for GPs aim to recover a target function with GPs in general, BIS recovers the target density with the design points guided by the GP bandits. We establish convergence guarantees for the weighted samples of BIS, independently of the convergence of the GP surrogate model itself.
Space-filling sequences—such as the Halton sequence (Halton, 1964)—can serve well as the candidate pool used in BIS, from which the samples are selected. It is a celebrated result in quasi-Monte Carlo (QMC; Dick and Pillichshammer, 2010) that numerical integration based on $N$ points from a space-filling sequence achieves the convergence rate $\mathcal{O}((\log N)^{d-1}N^{-1})$. This is faster than the concentration rate of Monte Carlo integration, $\mathcal{O}(N^{-1/2})$, for a large sample size $N$ (Morokoff and Caflisch, 1994). Yet, importance sampling with the space-filling sequence can be inefficient for small to medium sample sizes $N$. This is because space-filling sequences are not adaptive to the high-probability region of the target density, as they are designed to explore the domain uniformly. Using a space-filling sequence in BIS corresponds to changing the order of the space-filling sequence so that the high-probability region of the target density is explored first. This allows for efficient approximation accuracy for both small and large sample sizes $N$.
@@
The rest of the paper is structured as follows. Section˜2 concisely recaps importance sampling and GP regression. Section˜3 introduces the methodology of BIS in full. Section˜4 establishes a theoretical analysis of the rate of the convergence. Section˜5 examines the empirical performance of BIS, with applications to Bayesian inference of computationally-expensive models. Relevant work on GP-based experimental designs for Bayesian computation is recapped in Section˜6. Finally, Section˜7 concludes the paper with a discussion of future directions.
@@ 2 Background
This section provides a concise overview of importance sampling and GP regression. Let $\Theta\subset\mathbb{R}^{d}$ for some dimension $d$. Any probability distribution on $\Theta$ considered in this work is a Borel probability distribution and identified with the density with respect to the Lebesgue measure if it exists.
@@ 2.1 Importance Sampling
Importance sampling is one of the most fundamental sampling methods, which is suitable when drawing samples from a density of interest $p(\theta)=q(\theta)/Z$ on $\Theta$ is undesirable. The aim is to produce a set of $M$ weighted samples, $\{\theta_{n},w_{n}\}_{n=1}^{M}$, which approximates the density $p(\theta)$. The samples $\{\theta_{n}\}_{n=1}^{M}$ are drawn from a proposal density $u(\theta)$ in lieu of $p(\theta)$. If the density $p(\theta)$ is known in full, the weights $\{w_{n}\}_{n=1}^{M}$ are each set to the ratio $p(\theta_{n})/u(\theta_{n})$. If the density $p(\theta)$ is known only up to the normalisation constant $Z$, the self-normalised weights $\{w_{n}\}_{n=1}^{M}$ are used, where the expectation of an integrand $f(\theta)$ with respect to $p(\theta)$ is estimated as
$\mathbb{E}_{\theta\sim p}[f(\theta)]\approx\sum_{i=1}^{M}w_{n}f(\theta_{n})\quad\text{where}\quad w_{n}=\frac{q(\theta_{n})/u(\theta_{n})}{\sum_{n=1}^{M}q(\theta_{n})/u(\theta_{n})}.$ (1)
@@
The weights depend on the proportional $q(\theta)$ only, as the self-normalisation cancels out the constant $Z$. Consistency of the Monte Carlo estimate above is well established.
For importance sampling to be efficient, a sufficient portion of the samples $\{\theta_{n}\}_{n=1}^{M}$ should be situated well around the high-probability region of the target density $p(\theta)$. Otherwise, samples with insignificant weights near zero will be dominant, resulting in little contribution to the approximation of the density $p(\theta)$. The panel (a) in Figure˜1 illustrates a case where only around 10% of samples have significant weights while the other 90% are negligible. AIS improves the efficiency of importance sampling by optimising the proposal density $u(\theta)$ to draw samples. In AIS, a parametric family of proposal densities is typically specified, and samples are drawn from the family after or along the optimisation process. This approach is, however, not ideal for our setting, as successful optimisation of the proposal density $u(\theta)$ requires intensive evaluations of the target density $p(\theta)$. On top of that, as the samples are random and independent, some can be redundant, failing to be the most informative point to evaluate the target density $p(\theta)$.
@@
Suppose that the density $p(\theta)$ is so computationally expensive that the proportional $q(\theta)$ can be evaluated only up to $N$ times. A natural desideratum under such circumstances is to find $N$ samples that lead to the most accurate approximation of the density $p(\theta)$ via importance sampling. In the panel (a) in Figure˜1, for example, where 200 samples are drawn, selecting $N$ samples among them with the largest weights achieves the best accuracy under $N$ density evaluations. To this end, we cast sampling as a bandit problem to sequentially select $N$ most effective points from a pool of $M$ well-dispersed points that can be easily drawn. See the panel (b) in Figure˜1 for illustration.
@@
Figure 1: Illustration of two sampling schemes (a) and (b) applied to a normal density (solid line). The left panel shows 200 quasi-uniform points used as samples, where the points associated with the top 10% weights are circular, and the rest are triangular. The right panel shows 30 samples selected by BIS from the same 200 quasi-uniform points, where BIS exhausted all the points with the top 10% weights within 30 density evaluations.
@@ 2.2 GP Regression
A GP, denoted by $\mathcal{GP}(m,k)$, is a stochastic process determined by a mean function $m:\Theta\to\mathbb{R}$ and a covariance kernel $k:\Theta\times\Theta\to\mathbb{R}$. It can be understood as a random variable $f\sim\mathcal{GP}(m,k)$ taking values in a space of functions on $\Theta$. A characteristic property of GPs is the moment conditions
$m(\theta)=\mathbb{E}_{f\sim\mathcal{GP}(m,k)}[f(\theta)]\quad\text{and}\quad k(\theta,\theta^{\prime})=\mathbb{E}_{f\sim\mathcal{GP}(m,k)}[(f(\theta)-m(\theta))(f(\theta)-m(\theta^{\prime}))],$ (2)
which hold pointwise. Another characteristic property of GPs is that the joint law of the image $(f(\theta_{1}),\dots,f(\theta_{n}))$ of inputs $(\theta_{1},\dots,\theta_{n})$, under the sample path $f\sim\mathcal{GP}(m,k)$, follows a multivariate Gaussian distribution $\mathcal{N}(\bm{m}_{n},\bm{K}_{n})$. Here the mean $\bm{m}_{n}\in\mathbb{R}^{n}$ is a vector whose $i$-th element is $m(\theta_{i})$ and the covariance $\bm{K}_{n}\in\mathbb{R}^{n\times n}$ is a matrix whose $(i,j)$-th element is $k(\theta_{i},\theta_{j})$. Common choices of the covariance kernel $k$ include Gaussian and Matérn kernels. See e.g. Rasmussen and Williams (2005) for a more detailed introduction to GPs.
GPs provide a framework for nonparametric Bayesian modelling of regression problems. Suppose that a set of inputs $(\theta_{1},\dots,\theta_{n})$ and the associated outputs $(y_{1},\dots,y_{n})$ are observed as data. Given a GP prior distribution $\mathcal{GP}(m,k)$ and the observed data $(\theta_{i},y_{i})_{i=1}^{n}$, the posterior distribution remains as a GP, denoted by $\mathcal{GP}(m_{n},k_{n})$, where the mean function $m_{n}$ and covariance kernel $k_{n}$ are updated from those of the GP prior as follows:
$m_{n}(\theta)$ $=\bm{k}_{n}(\theta)^{\mathrm{T}}\bm{K}_{n}^{-1}(\bm{y}_{n}-\bm{m}_{n}),$ (3)
@@
Here, $\bm{k}_{n}(\theta)$ denotes a vector-valued function defined by $\bm{k}_{n}(\theta)=(k(\theta,\theta_{1}),\cdots,k(\theta,\theta_{n}))\in\mathbb{R}^{n}$, and $\bm{y}_{n}\in\mathbb{R}^{n}$ denotes a vector whose $i$-th element is $y_{i}$. In practice, it is common to assume that each output value is realised under i.i.d. observation noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$, in which case $\bm{K}_{n}$ is replaced with $\bm{K}_{n}+\sigma\bm{I}_{n\times n}$—where $\bm{I}_{n\times n}$ is the $n\times n$ identity matrix—in the above definition.
3 Methodology
We now introduce BIS, a method that frames the design of adaptive samples as a bandit problem. Section˜3.1 presents a general framework of BIS. It is built on two components: (i) a pool of candidates from which samples are selected; (ii) a criterion to select a sample at each iteration. Section˜3.2 discusses our default choice of the candidate pool. Section˜3.3 defines our point-selection criterion, termed GP upper Jensen bound (GP-UJB), which capitalises on a GP surrogate density. Section˜3.4 provides a discussion of the choice of hyperparameters involved in GP-UJB.

### v2

@@ Abstract
We propose bandit importance sampling (BIS), a powerful importance sampling framework tailored for settings in which evaluating the target density is computationally expensive. BIS facilitates accurate sampling while minimizing the required number of target-density evaluations. In contrast to adaptive importance sampling, which optimizes a proposal distribution, BIS directly optimizes the set of samples through a sequential selection process driven by multi-armed bandits. BIS serves as a general framework that accommodates user-defined bandit strategies. Theoretically, the weak convergence of the weighted samples, and thus the consistency of the Monte Carlo estimator, is established regardless of the specific strategy employed. In this paper, we present a practical strategy that leverages Gaussian process surrogates to guide sample selection, adapting the principles of Bayesian optimization for sampling. Comprehensive numerical studies demonstrate the superior performance of BIS across multimodal, heavy-tailed distributions, and real-world Bayesian inference tasks involving Markov random fields.
@@ 1 Introduction
Statistical characterization of complex phenomena often necessitates highly structured, black-box models that are computationally expensive to evaluate (Chen2010). While Bayesian inference provides a coherent probabilistic framework for integrating observations with prior knowledge, its application to black-box models is often computationally impractical. Such challenges are prevalent in fields, such as computational physics, biology, and geostatistics, where likelihood evaluations involve numerical solutions to dynamical equations (Beaumont2002; Girolami2008; Marin2012; Warne2020) or operations on high-dimensional matrices and massive summations (Rue2005; Goodreau2009; Besag1986).
@@
A fundamental task in Bayesian inference is to sample from a posterior density of the model parameter, which is proportional to the likelihood and the prior density. Conventional sampling methods, such as Markov chain Monte Carlo (MCMC; Brooks2011), necessitate extensive evaluations of the target density to explore the parameter space. The cumulative computational cost easily becomes prohibitive when the target density is expensive to evaluate. Consequently, there is a pressing need for evaluation-efficient sampling methods tailored for black-box densities, which can operate within a limited computational budget.
@@
Importance sampling—originally introduced as a variance reduction technique in Monte Carlo estimation (Hammersley1954)—offers a compelling approach in this context, approximating the target density by a weighted average of samples drawn from an arbitrary proposal density (Tokdar2010). Since target-density evaluations are strictly limited to the preselected samples, this approach allows for precise control over the computational budget. However, the approximation accuracy of importance sampling is heavily contingent upon the design of the samples. Adaptive importance sampling (AIS) is arguably the most prevalent approach to this challenge, optimizing the proposal density from which samples are drawn (e.g. Bugallo2017; Cotter2019; Cotter2020). Yet, AIS remains impractical for black-box densities, as the optimization of the proposal density reintroduces the need for intensive target-density evaluations.
In machine learning, Bayesian optimization (BO; Mockus1989) has emerged as a prominent framework for optimizing black-box objective functions that are expensive to evaluate (Shahriari2016). The core principle involves approximating the underlying objective function with a computationally efficient surrogate built upon the history of observed evaluations. Gaussian processes (GPs; Rasmussen2005) are a standard and versatile choice for surrogate modeling (Gramacy2020). GPs quantify predictive uncertainty alongside predictions, accounting for the epistemic uncertainty inherent in estimating the true function from finite data. BO iteratively selects the query point for evaluation by maximizing an acquisition function constructed from the GP surrogate model. This framework, sometimes referred to as GP bandits, can be cast as a bandit problem with an infinite number of arms (Bull2011). Accordingly, this framework has been extensively leveraged for experimental design and active learning tasks (Srinivas2010).
@@
This work proposes bandit importance sampling (BIS), a novel importance sampling framework driven by multi-armed bandits. BIS directly optimizes the samples, sequentially selecting them from a candidate pool via a bandit strategy. This approach preserves the fundamental advantage of importance sampling, limiting the number of target-density evaluations strictly to the pre-specified sample size. BIS is a general framework compatible with arbitrary bandit strategies. Theoretically, we establish the weak convergence of the weighted samples to the target density independently of the employed strategy. We present a practical implementation of the bandit strategy utilizing GP surrogate models, translating the principles of BO into sampling tasks. We demonstrate the efficiency of BIS across diverse experiments, including sampling from the marginal posterior of a Markov random field (MRF) model applied to large-scale precipitation data in the United States.
Scope
In this work, we restrict our attention to computationally expensive densities on low-dimensional spaces. The scalability and accuracy limitations of GPs in high-dimensional spaces remain open challenges. Sampling from high-dimensional densities poses its own set of challenges. The compound challenge of computational intensity and high dimensionality renders the problem intractable. Therefore, we isolate the challenge of computational intensity. Crucially, this setting retains substantial practical relevance. For example, in the Lorenz weather forecast model (Wilks2005), the posterior evaluation involves 40-dimensional stochastic dynamics, while the parameter space is two-dimensional. In hierarchical MRF models for massive spatial data (Rue2005), the high-dimensional latent field can often be marginalized out analytically, reducing the inference task to a few hyperparameters, yet the marginal posterior remains computationally expensive.
@@
The rest of the paper is structured as follows. Section˜2 recaps importance sampling and GP regression concisely. Section˜3 introduces the general framework of BIS. Section˜4 establishes a theoretical analysis of the convergence rate of BIS. Section˜5 develops a practical implementation of the bandit strategy for BIS, which leverages GP surrogate modeling. Section˜6 examines the empirical performance of BIS, with applications to Bayesian inference of computationally-expensive models. Relevant work on GP-based experimental designs for Bayesian computation is recapped in Section˜7. Finally, Section˜8 concludes the paper with a discussion of future directions.
@@ 2 Background
This section provides a concise overview of importance sampling and GP regression. Let $\Theta\subset\mathbb{R}^{d}$ for some dimension $d$. Any probability distribution on $\Theta$ considered in this work is identified with the density with respect to the Lebesgue measure if it exists.
@@ 2.1 Importance Sampling
Importance sampling is a fundamental technique, particularly effective when direct sampling from a target density $p(\theta)=q(\theta)/Z$ on $\Theta$ is impractical. The objective is to construct a set of $M$ weighted samples, $\{\theta_{n},w_{n}\}_{n=1}^{M}$, which approximates the target density $p(\theta)$. The samples $\{\theta_{n}\}_{n=1}^{M}$ are drawn from a proposal density $u(\theta)$ instead of $p(\theta)$. If the density $p(\theta)$ is known exactly, the importance weights $\{w_{n}\}_{n=1}^{M}$ are computed as the ratio $p(\theta_{n})/u(\theta_{n})$. If the density $p(\theta)$ is known only up to the normalization constant $Z$, the self-normalized weights $\{w_{n}\}_{n=1}^{M}$ are employed, by which the expectation of a test function $f(\theta)$ with respect to $p(\theta)$ is estimated as
$\mathbb{E}_{\theta\sim p}[f(\theta)]\approx\sum_{n=1}^{M}w_{n}f(\theta_{n})\quad\text{where}\quad w_{n}=\frac{q(\theta_{n})/u(\theta_{n})}{\sum_{n=1}^{M}q(\theta_{n})/u(\theta_{n})}.$ (2)
@@
The weights depend solely on the unnormalized target $q(\theta)$, as the self-normalization cancels the constant $Z$. Consistency of this Monte Carlo estimator is well established.
For importance sampling to be efficient, a substantial portion of the samples $\{\theta_{n}\}_{n=1}^{M}$ must be located within the high-probability regions of the target density $p(\theta)$. Otherwise, the vast majority of samples will be assigned negligible weights, contributing minimally to the representation of the target density $p(\theta)$. Panel (a) in Figure˜1 illustrates a scenario where only approximately 10% of the samples carry significant weights, while the remaining 90% are negligible. AIS enhances the efficiency of importance sampling by optimizing the proposal density $u(\theta)$ to draw samples. In AIS, a parametric family of proposal densities is typically employed, and samples are drawn from the family after or during the optimization process. However, this approach is impractical for our setting, as the successful optimization of the proposal density $u(\theta)$ necessitates extensive evaluations of the target density $p(\theta)$. Furthermore, since the samples are random and independent, they are prone to redundancy and often fail to be the maximally informative locations for evaluating $p(\theta)$.
@@
Suppose that the evaluation of $p(\theta)$ is so computationally expensive that the unnormalized target $q(\theta)$ can be evaluated at most $N$ times. A natural objective given this constraint is to identify $N$ samples that yield the most accurate approximation of the density $p(\theta)$ via importance sampling. For instance, in Panel (a) in Figure˜1, where 200 samples are initially drawn, selecting the $N$ samples with the largest weights yields the best accuracy given a budget of $N$ density evaluations. To this end, we frame the problem as a sequential decision-making task to select the $N$ most informative points from a pool of $M$ well-dispersed candidates. See Panel (b) in Figure˜1 for illustration.
@@
Figure 1: Illustration of the sampling efficiency of BIS on a normal density (solid contour). Panel (a): Standard importance sampling with 200 quasi-uniform points; circles indicate the 20 largest weights (top 10%), and triangles indicate the rest. Panel (b): Selection by BIS with a budget of $N=30$ evaluations; the method identified all the high-weight points among the 200 candidates while evaluating the target only 30 times, achieving a 85% reduction in computational cost.
@@ 2.2 GP Regression
A GP, denoted by $\mathcal{GP}(m,k)$, is a stochastic process fully specified by its mean function $m:\Theta\to\mathbb{R}$ and covariance kernel $k:\Theta\times\Theta\to\mathbb{R}$. It can be interpreted as a random variable $f\sim\mathcal{GP}(m,k)$ taking values in a function space on $\Theta$. A defining characteristic of GPs is the moment condition:
$m(\theta)=\mathbb{E}_{f\sim\mathcal{GP}(m,k)}[f(\theta)]\quad\text{and}\quad k(\theta,\theta^{\prime})=\mathbb{E}_{f\sim\mathcal{GP}(m,k)}[(f(\theta)-m(\theta))(f(\theta)-m(\theta^{\prime}))].$ (3)
Another characteristic of GPs is that the joint distribution of the function values $(f(\theta_{1}),\dots,f(\theta_{n}))$ at any finite set of inputs $(\theta_{1},\dots,\theta_{n})$ follows a multivariate Gaussian distribution $\mathcal{N}(\bm{m}_{n},\bm{K}_{n})$. Here, the mean $\bm{m}_{n}\in\mathbb{R}^{n}$ denotes the vector with $i$-th entry $m(\theta_{i})$ and the covariance $\bm{K}_{n}\in\mathbb{R}^{n\times n}$ denotes the matrix with $(i,j)$-th entry $k(\theta_{i},\theta_{j})$. Widely-used covariance kernels include Gaussian and Matérn kernels. See, e.g., Rasmussen2005 for a detailed introduction.
GPs offer a flexible framework for nonparametric Bayesian regression. Suppose we observe a set of inputs $(\theta_{1},\dots,\theta_{n})$ and their corresponding outputs $(y_{1},\dots,y_{n})$. Conditioning the GP prior $\mathcal{GP}(m,k)$ on the observed data yields a posterior distribution that is also a GP, denoted by $\mathcal{GP}(m_{n},k_{n})$, with the updated mean and covariance functions given by
$m_{n}(\theta)$ $=m(\theta)+\bm{k}_{n}(\theta)^{\mathrm{T}}\bm{K}_{n}^{-1}(\bm{y}_{n}-\bm{m}_{n}),$ (4)
@@
Here, $\bm{k}_{n}(\theta)=(k(\theta,\theta_{1}),\cdots,k(\theta,\theta_{n}))\in\mathbb{R}^{n}$ is the vector of kernel evaluations between the test point $\theta$ and the observed inputs, and $\bm{y}_{n}=(y_{1},\dots,y_{n})\in\mathbb{R}^{n}$ is the vector of the observed outputs. In practice, it is common to assume that observations are subject to i.i.d. Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma_{*}^{2})$ with the scale $\sigma_{*}$. In this case, $\bm{K}_{n}$ is replaced with $\bm{K}_{n}+\sigma_{*}^{2}\bm{I}_{n\times n}$ in the aforementioned definition, where $\bm{I}_{n\times n}$ is the $n\times n$ identity matrix.
3 General Framework
We now introduce BIS, a method that frames the adaptive design of samples as a bandit problem. This section presents a general framework of BIS, which comprises two key components: (i) a pool of candidates from which samples are selected, and (ii) a criterion for selecting a sample at each iteration. Specifying these two components (i) and (ii) defines a concrete instance of BIS. The construction of an effective candidate pool for black-box densities is detailed in Section˜3.2. The construction of an effective point-selection criterion, which is driven by GP bandits in this work, will be elaborated in the subsequent section. The standing setup for BIS is stated below.